OpenStudy (anonymous):

Can someone walk me through the steps? What is the center and radius of a circle with the given equation? x^2+y2-4x+10y=7

OpenStudy (anonymous):

Equation of a circle is (x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

where (a,b) is the center of the circle.

OpenStudy (anonymous):

Right :)

OpenStudy (anonymous):

Alright, so x^2 - 4x _ y^2 + 10y = 7 You need to completes the squares and get it into proper form.

OpenStudy (anonymous):

x^2 -4x + 4 + y^2 + 10y + 25 = 7 + 4 + 25

OpenStudy (anonymous):

(x-2)^2 + (y+5)^2 = 36

OpenStudy (anonymous):

center is (2,5) and radius is 6

OpenStudy (anonymous):

Oh i got my center as (-2,5)

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/completing-square.html Tutorial on completing the square if you need a refresher on that.

OpenStudy (anonymous):

Sorry, center is (2, -5)

OpenStudy (anonymous):

(x-a) is when a is positive so if you have (x+a) then a must be negative because (x-(-a)) = (x+a)

OpenStudy (anonymous):

(x - (+a))^2 + (y-(+b)^2 = r^2 is the eq of circle

OpenStudy (anonymous):

if the + signs help you see that

OpenStudy (anonymous):

Got it, so the center is (2,-5) and the radius is 6 I'm such an idoit i got 36 duh.. I understand what i did wrong thank you :)

OpenStudy (anonymous):

It's a common mistake! All that matters is that you are trying and succeeding. :)